# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ 2. Preliminary results
$\lim_{R\to\infty}\lim_{k\to\infty}\int_{B^{c}_{R}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (6.24)
In virtue of (6.23) and (6.24), we deduce that (6.21) holds. Hence we obtain from (6.17), (6.20), Lemma 6.6 and the Lebesgue dominated convergence theorem that
$\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u\Delta\psi\,\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u\Delta\psi~{}\mathrm{d}x=\int_{\mathbb{R}^{n}}\frac{g(x,u)}{|x|^{\gamma}}\psi\,\mathrm{d}x,~{}\forall~{}\psi\in C^{\infty}_{0}(\mathbb{R}^{n}).$
On exploiting the density of $C^{\infty}_{0}(\mathbb{R}^{n})$ in $E$, it follows that $u$ is a weak solution of ($\boldsymbol{S}$).
To complete the proof, it is sufficient to show the compactness of the $(PS)_{c}$ sequence $\{u_{k}\}_{k}$ in $E$. Invoking the fact that $c>0$, we have two possibilities as discussed below.
Case-1:$(c\neq 0,~{}u=0)$ In this situation, from (6.21), one has
$\int_{\mathbb{R}^{n}}\frac{G(x,u_{k})}{|x|^{\gamma}}\to 0\quad\text{as}\quad k\to\infty.$
@@
The above inequality together with $c\in(0,c_{0})$ implies the validity of (6.4), we thank to Lemma 6.5. Hence, up to a subsequence, still not relabelled, we can assume that
@@
Let us fix $m\in(\|u_{k}\|^{\frac{n}{n-2}},\frac{\beta_{\gamma,n}}{\alpha_{0}})$ and choose $r$, $r^{\prime}$ such that $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. Pick $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $r>1$ close to $1$ such that $\alpha rm<\beta_{\gamma,n}$. Therefore, by using (1.1), Theorem 1.3, Corollary 2.6, Theorem 1.7 and Hölder’s inequality, we have
@@
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})u_{k}}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (6.25)
Hence we deduce from $\langle{J^{\prime}(u_{k}),u_{k}}\rangle\to 0$ as $k\to\infty$ and (6.25) that $u_{k}\to 0$ in $E$ as $k\to\infty$. Consequently, since $J\in C^{1}(E,\mathbb{R})$, one can easily obtain $c=0$, which is a contradiction. Thus this situation is impossible.
Case-2:$(c\neq 0,~{}u\neq 0)$ In view of $u_{k}\rightharpoonup u$ in $E$ as $k\to\infty$, we obtain from the lower semicontinuity of the norm that, up to a subsequence, still denoted by iteslf, such that $\|u\|\leq\lim_{k\to\infty}\|u_{k}\|$. Indeed, if $\|u\|=\lim_{k\to\infty}\|u_{k}\|$, then by using the result of Autuori-Pucci [[6], Corollary A.2], one has $u_{k}\to u$ in $E$ as $k\to\infty$. It follows that the sequence $\{u_{k}\}_{k}$ satisfies the $(PS)_{c}$ compactness condition and thus $u$ is a nontrivial mountain pass solution of ($\boldsymbol{S}$), thanks to $J\in C^{1}(E,\mathbb{R})$.
@@
In view of (6.10), it is standard to check that $\|\psi_{k}\|=1$, $\psi_{0}\neq 0$ and $\psi_{k}\rightharpoonup\psi_{0}$ in $E$ as $k\to\infty$. Moreover, if $\|\psi_{0}\|=1$, then again by using [6, Corollary A.2], we have $u_{k}\to u$ in $E$ as $k\to\infty$. This contradicts $\|u\|<\lim_{k\to\infty}\|u_{k}\|$ and the proof is finished.
@@
Let $r$, $r^{\prime}>1$ be such that $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. From the above inequality, we infer that there esists $m>0$, $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $r>1$ close to $1$ such that
@@
Hence, from (1.1), Hölder’s inequality, Theorem 1.4, Theorem 1.7, Corollary 2.6, we have for large $k$ that
@@
×$\quad\times\bigg{(}\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{qr^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{qr^{\prime}}}\bigg{(}\int_{\mathbb{R}^{n}}\frac{\Phi_{m,j_{0}}(\psi_{k})}{|x|^{\gamma}}\,\mathrm{d}x\bigg{)}^{\frac{1}{r}}$
@@
Due to convexity of the map $s\mapsto\frac{1}{a}|s|^{a}$ for all $a\geq 1$, the above convergences imply that
$\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\,\mathrm{d}x=o_{k}(1)\quad\text{as}\quad k\to\infty$ (6.26)
for $t\in\{p,\frac{n}{2}\}$. Recall the well known Simon’s inequality as follows
$|x-y|^{\wp}\leq\begin{cases}C_{\wp}\big{[}\big{(}|x|^{\wp-2}x-|y|^{\wp-2}y\big{)}(x-y)\big{]}^{\frac{\wp}{2}}\big{[}|x|^{\wp}+|y|^{\wp}\big{]}^{\frac{2-\wp}{2}},\quad\text{if}~{}1<\wp<2,\\ c_{\wp}\big{(}|x|^{\wp-2}x-|y|^{\wp-2}y\big{)}(x-y),\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~{}\wp\geq 2,\end{cases}$ (6.27)
for all $x,y\in\mathbb{R}^{d}$, where $d\geq 1$, $C_{\wp}$ and $c_{\wp}$ are positive constants depending only on Weierstrass-$\wp$. By using (6.26) and (6.27), we first obtain for any $t\in\{p,\frac{n}{2}\}\geq 2$ that
$\int_{\mathbb{R}^{n}}|\Delta u_{k}-\Delta u|^{t}\,\mathrm{d}x$ $\lesssim\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\,\mathrm{d}x=o_{k}(1)\quad\text{as}\quad k\to\infty.$
Now assume that $t=p\in(1,2)$. Then according to (6.26), (6.27) and Hölder’s inequality, one has
$\int_{\mathbb{R}^{n}}|\Delta u_{k}-\Delta u|^{t}\,\mathrm{d}x$ $\lesssim\int_{\mathbb{R}^{n}}\Big{(}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big{)}\Delta(u_{k}-u)\Big{)}^{\frac{t}{2}}\Big{(}|\Delta u_{k}|^{t}+|\Delta u|^{t}\Big{)}^{\frac{2-t}{2}}\,\mathrm{d}x$
@@
In conclusion, we deduce that $\Delta u_{k}\to\Delta u$ in $L^{t}(\mathbb{R}^{n})$ as $k\to\infty$ for any $t\in\{p,\frac{n}{2}\}$ and thus $u_{k}\to u$ in $E$ as $k\to\infty$. Consequently, it contradicts $\|u\|<\lim_{k\to\infty}\|u_{k}\|$ and hence the result follows. ∎
@@
Let $\lambda_{0}$ be defined as in Lemma 6.4 and the hypothesis $(g_{6})$ holds for all $\lambda\geq\lambda_{0}$. Then by employing Proposition 6.7 and Theorem 6.2, the functional $J$ has a nontrivial critical point $u\in E$, which is a weak solution to ($\boldsymbol{S}$). This completes the proof. ∎
@@ Acknowledgements
Deepak Kumar Mahanta expresses his deep appreciation for the DST INSPIRE Fellowship, reference number DST/INSPIRE/03/2019/000265, supported by the Government of India. Tuhina Mukherjee expresses gratitude for the assistance received from the CSIR-HRDG grant; sanction no. 25/0324/23/EMR-II. Abhishek Sarkar received funding from the DST-INSPIRE Grant DST/INSPIRE/04/2018/002208, funded by the Government of India.

### v3

@@ 2. Preliminary results
(8.22) $\lim_{R\to\infty}\lim_{k\to\infty}\int_{B^{c}_{R}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=0.$
In virtue of (8.21) and (8.22), we deduce that (8.19) holds. Hence, we obtain from (8.15), (8.18), Lemma 8.6 and the Lebesgue dominated convergence theorem that
$\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u\Delta\psi\,\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u\Delta\psi~\mathrm{d}x=\int_{\mathbb{R}^{n}}\frac{g(x,u)\psi}{|x|^{\gamma}}\,\mathrm{d}x,~\forall~\psi\in C^{\infty}_{0}(\mathbb{R}^{n}).$
By exploiting the density of $C^{\infty}_{0}(\mathbb{R}^{n})$ in $E$, it follows that $u$ is a weak solution of the problem ($\mathcal{P}$). Now, to complete the proof, it is sufficient to show the compactness of the $(PS)_{c}$ sequence $\{u_{k}\}_{k}$ in $E$. Invoking the fact that $c>0$, we have two possibilities as discussed below.
[…]
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{G(x,u_{k})}{|x|^{\gamma}}\,\mathrm{d}x=0.$
@@
The above inequality together with $c\in(0,c_{0})$ implies the validity of (8.4), we thank to Lemma 8.5. Hence, without loss of generality up to a subsequence, still not relabelled, we can assume that
@@
Let us fix $m\in(\|u_{k}\|^{\frac{n}{n-2}},\frac{\beta_{\gamma,n}}{\alpha_{0}})$ and choose $r$, $r^{\prime}>1$ such that $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. Pick $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $r>1$ close to $1$ such that $\alpha rm<\beta_{\gamma,n}$. Therefore, by using (1.14), Theorem 1.1, Corollary 2.6, Theorem 1.5 and the Hölder’s inequality, we have
@@
(8.23) $\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})u_{k}}{|x|^{\gamma}}\,\mathrm{d}x=0.$
Hence, we deduce from $\langle{J^{\prime}(u_{k}),u_{k}}\rangle\to 0$ as $k\to\infty$ and (8.23) that $u_{k}\to 0$ in $E$ as $k\to\infty$. Consequently, since $J\in C^{1}(E,\mathbb{R})$, one can easily obtain $c=0$, which is a contradiction. Thus, this situation is impossible.
[…]
@@
In view of (8.8), it is standard to check that $\|\psi_{k}\|=1$, $\psi_{0}\neq 0$ and $\psi_{k}\rightharpoonup\psi_{0}$ in $E$ as $k\to\infty$. Moreover, if $\|\psi_{0}\|=1$, then again by using [6, Corollary A.2], we have $u_{k}\to u$ in $E$ as $k\to\infty$. This contradicts $\|u\|<\lim_{k\to\infty}\|u_{k}\|$ and thus, the proof is finished.
@@
Let $r$, $r^{\prime}>1$ be such that $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. From the above inequality, we infer that there exists $m>0$, $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $r>1$ close to $1$ such that
@@
By using (1.14), the Hölder’s inequality, Theorem 1.2, Theorem 1.5, Corollary 2.6, we have for large $k$ that
@@
$\quad\times\bigg(\int_{\mathbb{R}^{n}}\frac{|u_{k}-u|^{qr^{\prime}}}{|x|^{\gamma}}\,\mathrm{d}x\bigg)^{\frac{1}{qr^{\prime}}}\bigg(\int_{\mathbb{R}^{n}}\frac{\Phi_{m,j_{0}}(\psi_{k})}{|x|^{\gamma}}\,\mathrm{d}x\bigg)^{\frac{1}{r}}$
@@
Due to convexity of the map $s\mapsto\frac{1}{\eta}|s|^{\eta}$ for all $\eta\in[1,+\infty)$, the above convergences imply that
(8.24) $\int_{\mathbb{R}^{n}}\big(|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big)\Delta(u_{k}-u)\,\mathrm{d}x=o_{k}(1)\quad\text{as}\quad k\to\infty$
for $t\in\{p,\frac{n}{2}\}$. Recall the well-known Simon’s inequality as follows
(8.25) $|x-y|^{\eta}\leq\begin{cases}C_{\eta}\big[\big(|x|^{\eta-2}x-|y|^{\eta-2}y\big)(x-y)\big]^{\frac{\eta}{2}}\big[|x|^{\eta}+|y|^{\eta}\big]^{\frac{2-\eta}{2}}\quad\text{if}~1<\eta<2,\\ c_{\eta}\big(|x|^{\eta-2}x-|y|^{\eta-2}y\big)(x-y)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~\eta\geq 2,\end{cases}$
for all $x,y\in\mathbb{R}^{d}$, where $d\in[1,+\infty)$, $C_{\eta}$ and $c_{\eta}$ are positive constants depending only on $\eta$. From (8.24) and (8.25), we obtain for all $t\in\{p,\frac{n}{2}\}\geq 2$ that
$\|\Delta u_{k}-\Delta u\|^{t}_{t}$ $\lesssim\int_{\mathbb{R}^{n}}\big(|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big)\Delta(u_{k}-u)\,\mathrm{d}x=o_{k}(1)\quad\text{as}\quad k\to\infty.$
Similarly, from (8.24), (8.25) and the Hölder’s inequality, one can deduce for $t=p\in(1,2)$ that
$\|\Delta u_{k}-\Delta u\|^{t}_{t}$ $\lesssim\int_{\mathbb{R}^{n}}\Big(\big(|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u\big)\Delta(u_{k}-u)\Big)^{\frac{t}{2}}\Big(|\Delta u_{k}|^{t}+|\Delta u|^{t}\Big)^{\frac{2-t}{2}}\,\mathrm{d}x$
@@
Hence, we deduce that $\Delta u_{k}\to\Delta u$ in $L^{t}(\mathbb{R}^{n})$ as $k\to\infty$ for any $t\in\{p,\frac{n}{2}\}$ and thus, we obtain $u_{k}\to u$ in $E$ as $k\to\infty$. Moreover, it contradicts $\|u\|<\lim_{k\to\infty}\|u_{k}\|$ and thus, the result follows. ∎
@@
Let $\lambda_{0}$ be defined as in Lemma 8.4 and the hypothesis ($\bf g{6}$) holds for all $\lambda\geq\lambda_{0}$. Now, by employing Proposition 8.7 and Theorem 8.2, one can see that the functional $J$ has a nontrivial critical point $u\in E$, which is a weak solution to the problem ($\mathcal{P}$). This completes the proof. ∎
@@ Acknowledgements
Deepak Kumar Mahanta expresses his deep appreciation for the DST INSPIRE Fellowship, reference number DST/INSPIRE/03/2019/000265, supported by the Government of India. Tuhina Mukherjee expresses gratitude for the assistance received from the CSIR-HRDG grant, sanction number 25/0324/23/EMR-II. Abhishek Sarkar received funding from the DST-INSPIRE Grant DST/INSPIRE/04/2018/002208, funded by the Government of India.